Fill in the blank to complete this equation:
1+2+3+...+(n-1)+n+n+(n-1)+...+3+2+1=___________
I know your suppose to square the (n-1), but I just don't know where to even start.

Answer by mananth(16946)

(1+2+3+...+(n-1)+n)+(n+(n-1)+...+3+2+1)=___________
It consists of 2 same sequences
Sum of n terms
Sn = n/2(2a+(n-1)d)
a=1
d=1
n=n

sn = n/2(2+(n-1)1))
Sn = n(n+1)/2
2 Sn = n(n+1)
n(n+1)